# SLIM: Sim-to-Real Legged Instructive Manipulation via Long-Horizon Visuomotor Learning

### v1

@@ VIII-C Evaluation Protocol
We use the Standard object spatial layout as shown in Figure 8 in a real-world scene Lobby. By doing so, it is easier to maintain repeatability across all methods and seeds. Later in subsequent experiments, we will evaluate the robot in other scenes or with other layouts. For task objects, we use cubes and baskets of different colors. More details on these physical objects are detailed in Appendix E-A.
For each method of each random seed, we roll out the robot’s policy for 20 episodes with varied initial scene states (relative spatial positioning and object colors etc.) according to the evaluation protocol. For the detailed protocol, please refer to Appendix E-B.
TABLE I: Real-world task success rates (mean±$\pm$±stddev), averaged across 3 random seeds, with 20 episodes per seed.
Method Cumulative Success Rate [%] $\uparrow$ Episode Time [s] $\downarrow$
Search+MoveTo Grasp Search+MoveTo(WObj) DropInto (Full Task)
@@ VIII-D Results and Analysis
We conduct experiments on the baseline and the proposed methods, and measure their success rates in the real world. The results are summarized in Table I.
We run this experiment with 3 different random seeds (for low-level, expert and agent training) and measure the performance in real, using the following two metrics to characterize both accuracy and efficiency:
@@
    Cumulative Subtask Success Rate: it is defined as the success rate upto the specified subtask starting from the beginning of the full task.
@@
    Episode Time: it is defined as the time completing the full task from the beginning. Since we applied a time limit of $t_{\rm max}\!=\!90$s in evaluation, for failed episodes, we set their episode time as $t_{\rm max}$.
Then we calculate the mean and standard deviation across the seeds as the final aggregated metric values and the results are summarized in Table I.
It can be observed from the Table I that No Arm Retract baseline performs well for the initial two subtasks, but for later subtasks, it does not perform well. Note that even though it can move to the basket with the object with a success rate of $28.3\%$, it can only drop the object into the basket with a success rate of $5.0\%$. The major reason is that without arm retract reward, the arm is typically at a pose that is not friendly for learning dropping into. No Perturb baseline has the lowest grasping success rate. Actually, without perturbations (object, arm mount and control), the policy is not robust both during learning and can be stuck during exploration. This is reflected in the low and full task success rate, indicating the importance of perturbations for learning. No Visual Aug’s performance on the initial Search subtask is lower than the that of above baselines. The main reason is that without visual augmentation, it can be distracted by other task-irrelevant background objects and can sometimes have a false positive in detecting the target. This implies that the visual augmentation is important for improving the robustness of the robot’s vision. Distillation Only baseline has an overall success rate that is higher than No Arm Retract and No Perturb, reconfirming the importance of arm-retract and perturbations. On the other hand, its performance gap with distillation-guided RL (SLIM which uses both distillation loss and RL loss) shows that distillation loss is not enough for achieving a consistently high performance.
The proposed method SLIM achieves the highest subtask and full task success rate among all the compared methods. Moreover, SLIM is also the most efficient, taking only 43.8s to accomplish the full task on average.
VIII-E Common Failures for SLIM
We also recorded common failures for the full SLIM method. When grasping, in one case, Go1 stopped too late, and kicked the cube far away. In the other case, the low level policy due to sim-to-real gap shook and moved backward, causing the cube to be just out of reach. Drop into often fails because of either early dropping right outside of the basket, or Go1 kicking the basket back when backing toward the basket, right before dropping. We believe most of these can be fixed, in fact, during development, we have seen some of our best policies to be more robust in cases similar to these failures, and achieve around 95% success.
@@ VIII-E Common Failures for SLIM
VIII-F Analysis of Simulation and Real Performance
We also summarize the performance of different baseline methods as well as SLIM in Figure 10 (averaged over three seeds). The Sim result for each method is obtained by evaluating the corresponding policy in simulation under a setting that is the same as its training setting (in distribution). The Real result denotes the full task success rate when evaluating the policy in the real-world, following the Standard spatial layout as shown in Figure 8. It should be noted that different methods have different abilities and behaviors after their policies are transferred to real. Methods performs well in sim (e.g. the No Visual Aug baseline) does not necessarily translate to a high-performing policy in real. Without arm retract (No Arm Retract), the exploration because more difficult because the camera view is usually restrictive after grasping an object. In addition, it also causes a large sim-to-real gap, as moving with an outreaching arm will cause arm shakiness due to factors such as backlash and arm’s elasticity, which are hard to model in simulation. While one might expect perturbations to complicate the learning process, their inclusion result in much higher performance even in simulation. In the absence of perturbations (No Perturb), the policy is susceptible to converging to suboptimal solutions due to a lack of exposure to a sufficient range of task parameter combinations, effectively lacking a proper curriculum. The proposed SLIM performs well in simulation and maintains the highest success rate when transferred to real.
@@ VIII-F Analysis of Simulation and Real Performance
Figure 10: Full Task Success Rates of Different Methods in Simulation and Real World.
VIII-G Real-world Performance Across Different Scenes
In addition to the scene used in Table I, here we further run real-world evaluations of SLIM for the following scene variations (c.f. Figure 9 left): Outdoor, Carpet-Room, Kitchen and Lobby. For the first three scenes, we use the standard spatial layout (c.f. Figure 8). For Lobby, we use the Cluttered spatial layout, which is out of the training setting in simulation (Standard layout based). The results are summarized in Figure 9.
VIII-H Policy Behavior Visualization
The behavior of the policy is shown in Figure 11. To make the visualization clear we did not add background objects in the simulation for this episode. The robot start with a randomized position and yaw within a region (Start phase). A red cube and a green cube as well as a red basket and and blue basket are used. Their positions are randomized around the Standard layout and their yaw angles are also randomized. Given an instruction (“Drop the yellow cube into the red basket.”), the robot will first search for the target graspable object (the yellow cube). Note that during the Search phase, the robot actively moves both its body and the arm in searching of the target object (c.f. Figure 11-Search). Once the target object is in view, the robot will move towards it (Figure 11-MoveTo). Once the robot moves close to the target object, it will start to grasp (Figure 11-Grasp). During this stage, the robot body remains stationary and the arm will move towards the target cube and grasp it. After picking up the target cube, the robot will search for the target basket while keeping the cube in gripper. As shown in Figure 11-SearchWithObj, the robot rotates around its current position in searching of the target basket (as indicated by the circular ee-trajectory). At the same time, the arm retracts from the initial out-reaching pose after grasping towards a neutral arm pose, as indicated by the shrinking radius of the circular ee-trajectory in Figure 11-SearchWithObj. Once the target basket is in view, the robot will move towards it (Figure 11-MoveToWithObj) and drop the cube in gripper into that basket (Figure 11-DropInto).
@@ A-A Subtask Boundaries, Rewards and Behavior Priors
    $\mathcal{S}$(Search): succeeds when the target object is in the view of the gripper-mounted RGB camera;
@@
    $\mathcal{S}$(SearchWObj): succeeds when the target object is in gripper and the target basket enters the view of the gripper-mounted RGB camera;
@@
  • stationary: both the forward and angular velocity of the command to the quadruped are kept as zero ($\mathbf{c}=0$). This encourages the robot to be stationary. It is used in manipulation related tasks (Grasp, MoveGripperToWObj, DropInto).
In practice, both priors are implemented in a similar way, by initializing the teacher’s actor network for those subtasks with these priors to be of zero mean and small std for the corresponding action dimensions (e.g. the forward velocity dimension for rotational prior and both the forward and angular velocity dimensions for stationary prior). The reason for this design is to accommodate the KL-based distillation step in student learning, by providing a valid action distribution.
@@ A-B Low-level Task and Reward
TABLE IV: Environment Randomization Parameters
@@ B-A Teacher High Level
Each policy net $\pi^{k}$ (for subtask $k$) is implemented as as a 3 layer MLP, with the layer size as $512$ ($[512,512,512]$). The critic net for each subtask uses the same MLP structure and with the size of ($[512,512,512]$). Frame stacking is used with a stacking size of $5$.
The re-planning period is $5$, meaning the high-level policy run once every $5$ low-level steps.
We used four 3090 GPUs for training, with a batch size of $2000$, learning rate of $1{\rm e}^{-4}$, gradient clipping of $0.1$, number of parallel environment of $120$, and training iterations of $4$M.
@@ B-B Student High Level
Both the critic and the actor have three hidden layers of $(512,512,512)$. The actor outputs a squashed Gaussian distribution as in SAC [18]. Note that the RL and distillation gradients can be propagated all the way back to the policy input encoder and thus affect the representation learning.
Appendix C Sim2Real Techniques
C-A Dynamics Gap Techniques
@@
C-B Visual Gap Techniques
@@
Visual sim-to-real, especially for colors, is notoriously difficult [40]. We simplify the problem by only using four distinct colors in our experiments, namely, red, green (more like cyan), blue and yellow. These colors are sampled from real objects in different lighting conditions and are randomized in the HSV space during simulation training. Color samples are shown in Figure A.12.
@@
Appendix D Robot Implementation Details
D-A Control Latency Management

### v2

@@ VIII-C Evaluation Protocol
We use the Standard object spatial layout as shown in Figure 8 in a real-world scene Lobby. By doing so, it is easier to maintain repeatability across all methods and seeds. Besides the Standard in lobby, we also evaluate the robot in several other scenes with other layout. For task objects, we use cubes and baskets of different colors. More details on these physical objects are detailed in Appendix F-A.
For each method of each random seed, we roll out the robot’s policy for 20 episodes with varied initial scene setup (i.e. different positioning and colors of objects) according to the evaluation protocol. For the detailed protocol, please refer to Appendix F-B.
TABLE I: Real-world task success rates and episode time (mean±$\pm$±stddev over 3 random seeds), with 20 episodes per seed.
Method Fully Cumulative Success Rate [%] $\uparrow$ Episode Time [s] $\downarrow$
Autonomous Search+MoveTo Grasp Search+MoveTo(WObj) DropInto (Full Task)
@@ VIII-D Results and Analysis
We conduct experiments on the baselines and the proposed method, and measure their success rates in the real world. The results are summarized in Table I.
We run the whole training pipeline with three different random seeds to train low-level, the teacher and the student, and measure the performance of the final student policy in real, in terms of accuracy and efficiency:
@@
    Cumulative Subtask Success Rate: the success rate up to the specified subtask from the beginning of the episode.
@@
    Episode Time: the time spent completing the full task from the beginning. We apply a time limit of $t_{\rm max}\!=\!90$s. For failed episodes, we use $t_{\rm max}$ as episode time.
Finally, we report the mean and standard deviation across the seeds as the final aggregated metric value in Table I.
Table I shows how important each of the interventions used contribute to the final successes. No Arm Retract baseline performs well for the initial two subtasks of Search and MoveTo, but not for the later subtasks. Even though half of the times the robot can move to the basket carrying the cube (a cumulative success rate of $28.3\%$), it only successfully drops the object into the basket less than one fifth of the cases (a cumulative success rate of $5.0\%$). The main reason is that without arm retract reward, the arm could be at an over-extended pose, getting closer to the physical torque limits of the shoulder motors responsible for raising the arm. This leads to more arm shake, cube slipping out of the fingers, motor failures, and makes it harder for the arm to raise up to search for the basket and to hover above the basket to complete DropInto. No Perturb baseline has the lowest grasping success rate. Without perturbations (object, arm mount and control), the policy is not robust enough to deal with the variations in the location of the object relative to the robot. This is likely because during simulation training, the policy can get very efficient at approaching the target object and stopping just at the success boundary of the MoveTo subtask, so that the Grasp subtask usually starts with very similar locations of the target object relative to the robot, while in the real world such clear cut positioning is rarely seen. No Visual Aug’s performance on the initial Search subtask is lower than the that of the above baselines. The main reason is that without visual augmentations, it can be distracted by background objects irrelevant to the task, i.e. false positive detection of the target. This shows that visual augmentations are important for improving the robustness of the robot’s vision module. Distillation Only baseline simply imitates from the highly successful teacher policy, and can indeed achieve a relatively high success rate. However, its performance is not consistent across random seeds. Our full method (SLIM) uses distillation-guided RL, and results in consistently high success rates across seeds. It shows that distillation loss alone is not enough for robust policy learning. Additionally, task completion is also slower than the full method which additionally uses RL loss to optimize the policy.
Our full method SLIM achieves the highest subtask and full task success rate among all the methods, and is also the most efficient, taking only 43.8 seconds to complete the full task on average.
@@
VIII-E Generalization to Different Real-world Scenes
In addition to the standard scene used in Table I, we further run real-world evaluations of SLIM under more scene variations (c.f. Figure 9 left): Outdoor, Carpet-Room, Kitchen and Lobby. For the first three scenes, we use the standard spatial layout (c.f. Figure 8). For Lobby, we use the Cluttered spatial layout, which is different from the training (Standard) layout. Results as summarized in Figure 9 are close to the Standard layout or a bit lower due to being out of distribution. This shows that the model trained fully in simulation can zero-shot adapt to a wide range of real world settings, with vastly different lighting conditions, background, and floor materials and textures.
Furthermore, our qualitative evaluation showed that the robot also generalizes well to scenarios where:
  1. 1.
    task objects are randomly scattered on the ground (e.g., last two rows in Figure 1),
  2. 2.
    distractor objects are present (e.g., last row in Figure 1),
  3. 3.
    human interrupts the task progress, or
  4. 4.
    objects to grasp have slightly different shapes.
We refer the reader to our demo videos for more emergent generalization behaviors. In general, the robot is quite robust to the real-world scene complexity.
VIII-F Comparison with Simulation Performance
Different methods or even the same method with different seeds can exhibit very different behaviors, mostly due to the variability in the RL training of the teacher policies. Additionally, because of the sim-to-real gap, real deployment can differ from simulation as well.
Figure 10 shows simulation performance with the real performance for our methods, so we know more clearly how each method helps the training process.
The Sim success rate for each method denotes the average full task success rate over 200 episodes run in simulation, averaged across the three random seeds. The simulation environment follows the same randomization setting as training, with the policy rollout being fully greedy. The Real success rate denotes the full task success rate of the policy in the real-world, following the Standard spatial layout as shown in Figure 8.
Without arm retract (No Arm Retract), searching for the basket becomes more difficult in simulation as well as in real, because after grasp, the camera still points down to the ground, but searching for the basket requires many steps of the arm raising up, creating an exploration bottleneck for RL training. Real world performance suffers more, because moving with an out-stretched arm causes large sim-to-real gaps: The tip of the arm tends to shake more due to motor backlash, hard to model in simulation, and the shoulder motors operate close to their physical torque limits, leading to more motor overload. For No Perturb, while one might expect perturbations to complicate RL learning, their inclusion results in a much higher performance even in simulation. Without perturbations, the policy seems to converge to suboptimal solutions maybe due to a lack of exposure to a sufficiently large range of setups. Without visual augmentations (No Visual Aug), simulation task is easier to solve, thus simulation success is artificially high, but real world performance suffers when compared to the full method. RL training loss in the full SLIM method increases success in simulation (over the imitation only Distillation Only) and achieves the highest success rate in real. It’s interesting to note that the RL loss seems to help the No Visual Aug baseline to achieve slightly better real world success than Distillation Only.
@@ VIII-F Comparison with Simulation Performance
Figure 10: Full Task Success Rates of Different Methods (averaged across three seeds) in Simulation and Real.
@@
VIII-G Policy Behavior Visualization
The behavior of the policy is shown in Figure 11. To make the visualization clear we did not add background objects in the simulation for this episode. The robot start with a randomized position and yaw within a region (Start phase). A red cube and a green cube as well as a red basket and and blue basket are used. Their positions are randomized around the Standard layout and their yaw angles are also randomized. Given an instruction (“Drop the yellow cube into the red basket.”), the robot will first search for the target graspable object (the yellow cube). Note that during the Search phase, the robot actively moves both its body and the arm in searching of the target object (c.f. Figure 11-Search). Once the target object is in view, the robot will move towards it (Figure 11-MoveTo). Once the robot moves close to the target object, it will start to grasp (Figure 11-Grasp). During this stage, the robot body remains stationary and the arm will move towards the target cube and grasp it. After picking up the target cube, the robot will search for the target basket while keeping the cube in gripper. As shown in Figure 11-SearchWithObj, the robot rotates around its current position in searching of the target basket (as indicated by the circular ee-trajectory). At the same time, the arm retracts from the initial out-reaching pose after grasping towards a neutral arm pose, as indicated by the shrinking radius of the circular EE trajectory in Figure 11-SearchWithObj. Once the target basket is in view, the robot will move towards it (Figure 11-MoveToWithObj) and drop the cube in gripper into that basket (Figure 11-DropInto).
VIII-H Failure Modes
We also recorded common failure cases for SLIM. In one case, when grasping the quadruped stopped too late and kicked the cube far away. In the other case, the low-level policy due to sim-to-real gap shook and moved backward, causing the cube to be just out of reach. DropInto sometimes fails because of either early dropping right outside of the basket, or the quadruped kicking the basket away right before dropping when backing towards the basket. In rare cases where the gripper holds only the upper half of a cube, the robot tends to get stuck at hovering over the basket while holding the cube, because the cube is completely blocked by the gripper fingers in the wrist camera view. See our failure demo videos for more details.
@@ A-A Subtask Boundaries, Rewards and Behavior Priors
    $\mathcal{S}$(Search): succeeds when the target object is in the view of the wrist-mounted RGB camera;
@@
    $\mathcal{S}$(SearchWObj): succeeds when the target object is in gripper and the target basket enters the view of the wrist-mounted RGB camera;
@@
  • stationary: both the forward and angular velocity of the command to the quadruped are hardcoded as zero ($\mathbf{c}=0$). This encourages the robot to be stationary. It is used in manipulation related tasks (Grasp, MoveGripperToWObj, DropInto).
In practice, both priors are implemented in a similar way, by initializing the teacher’s actor network for those subtasks with these priors to be of zero mean and small std for the corresponding action dimensions (e.g. the forward velocity dimension for rotational prior and both the forward and angular velocity dimensions for stationary prior). The reason for this design is to accommodate the KL-based distillation step in student learning, by providing a valid action distribution and being consistent with subtasks without these priors.
@@ A-B Low-level Task and Reward
TABLE IV: Low-level Domain Randomization Parameters
@@ B-A Teacher High Level
Each policy net $\pi^{k}$ (for subtask $k$) is implemented as as a 3 layer MLP, with the layer size as $512$ ($[512,512,512]$). A squashed Normal projection layer (a layer for generating Normal distributions) is used as the output layer for each policy net following the standard practice [19], generating the action distribution conditioned on the current input.
The critic net for each subtask uses the same MLP structure and with the size of ($[512,512,512]$). Frame stacking is used with a stacking size of $5$.
The re-planning period is $5$, meaning that the high-level policy predicts once every $5$ low-level steps.
@@ B-B Student High Level
Both the critic and the actor have three hidden layers of $(512,512,512)$. The actor outputs a squashed Gaussian distribution as in SAC [19]. Note that the RL and distillation gradients will be propagated all the way back to the policy input encoder for learning the representation.
Stationary Bit.
The teacher policy incorporates a stationary behavior prior into the locomotion command $\mathbf{c}$ for several subtasks (Table II). With this prior, the corresponding locomotion command distribution always has a zero mean and a small std. Taking argmax of the distribution, the teacher policy generates exactly zero locomotion command and thus the quadruped can remain stationary. However, if we directly distill this distribution into the student’s counterpart, it is almost impossible to guarantee a zero mode, due to the continuous nature of the distribution. Thus we employed a sationary bit trick to solve this issue. Specifically, we create an extra binary action distribution for the student for predicting whether the quadruped should be stationary or not, which is determined by checking if the locomotion command of the teacher is exactly zero. We then add a new prediction loss for learning this stationary bit, while masking out the KLD loss (only for the two locomotion command dimensions) if the stationary bit is true. For student inference, whenever the stationary bit distribution outputs 1, the locomotion command $\mathbf{c}$ will be overwritten to be zero.
B-C Low Level
For the low-level training, we used PPO [44] with a regularized online adaptation module Fu et al. [15].
Policy Input and Locomotion Command.
As shown in Figure 2, the low-level locomotion policy takes two types of inputs: 1) the standard proprioceptive state $\mathbf{s}_{\textrm{leg}}$ obtained from the environment, 2) the locomotion command $\mathbf{c}$ from an external command issuer (e.g.. The second type of input is crucial to make the low-level policy useful in downstream tasks where locomotion is involved as a basic skill for exploration and learning. During training, the command is re-sampled periodically within each episode and the low-level policy is trained to follow the command $\mathbf{c}$ via the command-following rewards detailed in Table III. The arm pose is also adjusted periodically within each episode, to make the low-level policy’s command following robust against arm movement. For the standard proprioceptive input to the low-level policy, while it is also possible to include those from the arm. Empirically, we discover that it might trigger some shakiness behavior in certain cases and therefore we only used $\mathbf{s}_{\textrm{leg}}$.
After training, the low-level policy can be used to follow commands issued by an external command issuer, e.g. either a higher level policy (as in SLIM and all the other baseline except Human Teleop), or human tele-operator (as in Human Teleop) in this work.
Actor and Value Networks.
Both the actor and value networks have three hidden layers of size $(256,256,256)$. The actor network outputs a Beta distribution instead of the typical Gaussian. For numerical stability, we set the minimum value of the shape parameters of the Beta distribution to $1$.
Online Adaptation Module for Sim-to-Real Transfer
. Similar as in [15, Figure 2], our online adaptation module consists of two encoders, the privileged encoder and the adaptation encoder. The privileged encoder takes privileged information as input and predicts a latent representation of environment extrinsics, while the adaptation encoder predicts the same latent representation depending only on robot sensory observations. Both encoders are trained to match each other and jointly with the RL objective. The privileged encoder network has two hidden layers of $(128,64)$ and the adaptation encoder has hidden layers of $(256,128)$.
B-D Human Teleoperation
To show that we’ve made reasonable efforts to make Human Teleop work, we list its failure modes. 1) The most common failure mode is task timeout. This is sometimes due to not fully grasping the cube and the cube slipping out of the fingers during movement, leaving no time to search and re-grasp. Failing to grasp fully can be due to either stopping a bit far from the cube, or incorrectly assessing the distance to the cube under monocular vision. 2) The arm can enter a twisted pose due to poor IK solutions, despite having already tuned IK as we can.
Human teleoperation is in general less accurate and slower than RL policy, due to Low level policy shaking, and human operator fatigue. Besides being slower, human teleoperation seems to be less energy efficient than our RL policies. When RL learned high-level policies can run smoothly on a particular low level policy, human teleoperation on the same low level policy seems to trigger over-current protection of Go1. This often happens when trying to raise an over-extended arm, causing Go1 to lose control of its leg motors. In evaluations, we ignore these Go1 hardware failures, and let the human operator retry for free.
B-E Model Parameter Count
We summarize in Table V the rough number of model parameters for different components of our system, including low level, teacher high level and student high level.
TABLE V: Number of Trainable Parameters.
Component Approx. Param. Count
Low Level $3$M
Teacher High Level (all subtasks) $10$M
Student High Level $16$M
Our deployed policy (“Low Level” + “Student High Level”) has a total parameter count of less than 19M.
Appendix C Important Training Hyperparameters
Below we list some important training hyperparameters of SLIM.
Teacher High Level.
We used four 3090 GPUs for training, with a batch size of $2000$, a learning rate of $1\times 10^{-4}$, gradient clipping of $0.1$ with the Adam optimizer, $120$ parallel environments, and training iterations of $4$M (corresponding to about $60$M environment steps). The total training wall-clock time was about 1 week.
Student High Level.
We used 60 parallel environments to collect rollout data for training the student, where the data was stored in to a replay buffer with a capacity of $600$k environment steps. The full policy was trained on four 3090 GPUs for about 2 weeks, for a total of $1.3$M iterations. This amounts to roughly $40$M environment steps in total. We used a batch size of $64$, a learning rate of $1\times 10^{-4}$, gradient clipping of $100$, and the Adam optimizer. For the mixed rollout schedule, we kept the probability $\beta$ of sampling from the student policy as $0.5$ until the $750$K-th iteration, and linearly increased $\beta$ so that it reaches $1.0$ at the $1.25$M-th iteration. The distillation loss weight $\alpha$ in Equation 5 was set to $0.01$. The fixed modal dispersion $\sigma$ for distilling the teacher policy was $0.05$.
Low level.
We used 5000 parallel environments to roll out and train the low-level policy. The policy was trained on one 3090 GPU for about three days for a total of 4B environment steps. We used a batch size of 5000. The Adam optimizer had an initial learning rate of $2\times 10^{-4}$ and a step decay to its $1/5$ after $70\%$ of the total training progress.
Appendix D More Sim2Real Techniques
D-A More Dynamics Gap Reduction Techniques
@@
D-B More Vision Gap Reduction Techniques
@@
Vision sim-to-real, especially for colors, is notoriously difficult [41]. We simplify the problem by only using four distinct colors in our experiments, namely, red, green (more like cyan), blue and yellow. These colors are sampled from real objects in different lighting conditions and are randomized in the HSV space during simulation training. Color samples are shown in Figure A.12.
@@
Appendix E Robot Implementation Details
E-A Control Latency Management